OpenStudy (anonymous):

If a side-length of an equilateral triangle is 18, which is its area?

OpenStudy (anonymous):

do ure own work

OpenStudy (anonymous):

Sorry but Im a little confused the answers are. A.81 B.72 C.81sqrt(2 D.81sqrt(3

OpenStudy (2bornot2b):

\[1/2\times 6\times 6.sin(\frac{\pi}{3})\]

OpenStudy (2bornot2b):

Have you learnt trigonometry?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I come up with .3288

OpenStudy (2bornot2b):

I hope you know then that \[sin\frac{\pi}{3}=\frac{\sqrt3}{2}\]

OpenStudy (anonymous):

\[81\sqrt{3}\]

OpenStudy (farmdawgnation):

Austin Todd, you're more than welcome to try and draw more information out of the asker - in fact we encourage that. But whatever you do, please be polite. Thanks!

OpenStudy (anonymous):

I was just kidding with him, he's my bestfriend and we're both in the same geometry class, and we both needed help, and I'm actually the one who showed him this website. Ask him for yourself. I was just messing around with him.